# AutoTAMP: Autoregressive Task and Motion Planning with LLMs as Translators and Checkers

### v1

@@ 5.1 Task Success Rates
Table 2: Task success rates in multi-agent scenarios. Each scenario has hard temporal and geometric constraints.
@@
(multi-agent, hard time & geo. cst.)
GPT-3 end2end $0\%$ $0\%$ $0\%$
GPT-3 naive task planning $13.3\%$ $0\%$ $7\%$
GPT-3/STL $25\%$ $22\%$ $74\%$
GPT-3/STL/Syn. $70\%$ $35\%$ $85\%$
GPT-3/STL/Syn./Sem. (AutoTAMP) 89% 60.7% 89.7%
GPT-4 end2end $5\%$ $0\%$ $6\%$
GPT-4 naive task planning $17\%$ $0\%$ $47\%$
GPT-4/STL $85\%$ $46\%$ $95\%$
GPT-4/STL/Syn. $94\%$ $67\%$ $95\%$
GPT-4/STL/Syn./Sem. (AutoTAMP) 100% 79% 100%
@@
We report the task success rates for the single-agent and multi-agent scenarios in Table 1 and Table 2, respectively. For HouseWorld1 (Figure 4(a)) with no hard time constraint, we find that all methods using LLMs as task planners outperform our approach; this environment permits direct trajectories between any two positions and thus lacks geometric challenges that such methods will struggle with. When adding a strict time constraint (HouseWorld2), we see that such methods perform much worse while AutoTAMP’s success rate persists. For the remaining tasks that include geometric constraints, LLM End-to-end Motion Planning and Naive Task Planning both perform quite poorly. Unsurprisingly, we observe a general trend that GPT-4 outperforms GPT-3 for all methods.
5.2 Failure examples for LLM Task Planning and AutoTAMP Methods
LLM Task Planning: We observe that LLM Task Planning methods may fail at tasks that require long-horizon planning. Figure 5 below illustrates two examples. In (a), these methods inefficiently grab the key for a particular door only when that door is the next immediate obstacle. Rather than grab both key1 and key2 that are near the starting position before unlocking door1, it sequences the actions to grab key1, unlock door1, then grab key2. This inefficiency can lead to a violation of time constraints that are part of the task specification, resulting in task failure. In (b), ingredient1 is unnecessarily (and inefficiently) visited by both agents.
Figure 5: Two examples of failures by LLM Task Planning methods.
AutoTAMP: AutoTAMP (and the ablations) primarily fail because of translation errors. While the re-prompting techniques we use to improve translation are effective, there are still translation errors that result in task failure.
5.3 Planning Time
We report timing information for AutoTAMP with GPT-4 in Table 3. For each scenario, we report 10th, 50th, and 90th percentile runtime cost for the following steps in the overall process: GPT-4 API call, STL planner planning time, code execution time, GPT-4 translation to STL, syntactic check loop, semantic check loop. Overall, the STL planner runtime is the main bottleneck. In cases where the GPT-4 API call takes an excessive amount of time (i.e. more than 100s), we note that this is likely due to unstable latency and may not be a good proxy for decoding runtime.
We also report the number of iterations for both the syntactic and semantic re-prompting techniques for each task scenario. Note that semantic re-prompting subsequently requires an additional round of syntactic re-prompting for the new STL output; thus, syntactic re-prompting can have more total iterations than the limit of 5 since this limit is applied per STL output.
Table 3: Timing information for each step of AutoTAMP using GPT-4; Iteration counts for the syntactic and semantic re-prompting. For each, We report the 10th, 50th, and 90th percentile.
Time (s) HouseWorld2 Chip. Overcooked Rover Wall
Total Time $15/42/221$ $403/871/1579$ $34/77/157$ $79/123/174$ $21/43/67$
GPT-4 API $2/3/55$ $8/21/160$ $3/5/7$ $10/14/18$ $4/6/7$
STL Planner $10/35/152$ $310/732/1482$ $30/74/148$ $68/89/142$ $17/34/49$
Code Processing $/3/3$ 3 $3/4/4$ $/3/3$ 3 $3/4/4$ $/3/3$ 3
GPT-4/STL $5/9/10$ $172/457/681$ $16/20/45$ $15/18/29$ $6/9/15$
Syntactic Check $1/3/60$ $11/23/163$ $4/7/12$ $5/6/7$ $2/3/5$
Semantic Check $6/36/157$ $141/630/850$ $17/58/106$ $75/102/134$ $14/31/54$
Iterations HouseWorld2 Chip. Overcooked Rover Wall
Syntactic Check $1/2/2$ $3/5/7$ $1/2/2$ $2/3/6$ $1/2/3$
Semantic Check $1/2/3$ $/3/3$ 3 $2/2/3$ $2/3/3$ $2/2/3$
5.4 Ablation Studies
Syntactic and Semantic Checkers In Table 1 and 2, we also evaluate the impact of syntactic and semantic error correction on using LLMs to translate to STL. The results show that translation with no error correction has modest success across task scenarios, but both syntactic and semantic error correction significantly improve performance; this trend is present across all scenarios.
Fine-tuned NL2TL Translation Model Prior work (NL2TL [19]) has investigated translating from natural language to STL via a modular pipeline that utilizes a fine-tuned LLM. While incorporating NL2TL into our AutoTAMP framework was not part of our primary investigation, we report here its impact on task success in Table 4. The top three rows report AutoTAMP + NL2TL with ablations over syntactic and semantic re-prompting. The bottom three rows report the original ablation results of AutoTAMP without NL2TL. We see that integrating NL2TL generally provides a modest improvement. However, it is exciting that AutoTAMP appears competitive without the use of a fine-tuned model since it does not rely on a targeted dataset or additional offline training.
Table 4: Ablation studies of the AutoTAMP with/without integrating the finetuned NL2TL model.
HouseWorld2 Chip. Overcooked Rover Wall
GPT-4/STL/NL2TL $53\%$ $61\%$ $87\%$ $58\%$ $95\%$
GPT-4/STL/Syn./NL2TL $64\%$ $73\%$ $94\%$ $70\%$ $95\%$
GPT-4/STL/Syn./Sem./NL2TL 83.5% 86% 100% 79.7% 100%
GPT-4/STL $42\%$ $42.7\%$ $85\%$ $46\%$ $95\%$
GPT-4/STL/Syn. $59\%$ $70\%$ $94\%$ $67\%$ $95\%$
GPT-4/STL/Syn./Sem. 82% 87.7% 100% 79% 100%
6 Related Work
@@
LLMs for TAMP Recent claims about the impressive reasoning capabilities of LLMs [6, 24] have led to interest in such models for task and motion planning. One approach is to directly use LLMs as planners [7, 8, 9, 10, 11, 12, 13, 25]. Initial work showed that naive zero-shot generation of a primitive action sequence from a high-level task description had relatively poor executability, but few-shot in-context learning, constraining output to admissible actions, and autoregressive step-by-step action generation significantly improved performance [7]. Subsequent efforts connected to motion planning by grounding the primitive actions to motion control policies and using affordance functions to guide LLM-based task and motion planning [8, 9]; Huang et al. [10] closed the loop by prompting with feedback, such as action success indication, before querying for the next action. Other work focused on how prompting can inform task execution, such as few-shot summarization for user task preferences [11] or extraction of commonsense arrangements for underspecified object placement [13]. Despite these successes, however, there is some evidence that LLMs exhibit poor performance on more realistic tasks [15, 26], motivating different approaches. While we are interested in LLMs for TAMP, our work does not directly use LLMs as planners.
@@
Translating Language to Task Representations A natural alternative is to rely on dedicated planners by mapping from natural language to a planning representation. For over a decade, research has investigated mapping language to lambda calculus [27], motion planning constraints [28], linear temporal logic [29, 30, 31], and signal temporal logic [32, 33], among others. To address challenges of data availability, task generalization, and linguistic complexity, recent work has applied LLMs to this translation problem. Modular approaches have used LLMs to extract referring expressions with corresponding logic propositions to then construct a full temporal logic specification [34, 19]. Relying on LLMs for direct translation, other work has mapped from language to PDDL goals [17] or full PDDL problems [35, 16]. Our work similarly translates to a task specification, but we can represent complex constraints (e.g. temporal), and we introduce a novel mechanism for automatic detection and correction of semantic errors.
@@
Re-prompting of LLMs The quality of LLM output is greatly improved with useful context, such as few-shot in-context learning for novel tasks [6]. LLMs for TAMP are typically also provided task-relevant information, such as environment state or admissible actions [12]. Re-prompting with additional context based on LLM output has been shown to be extremely beneficial, such as with iterative action generation [7], environmental feedback [10], inadmissible actions [7, 8, 9], unmet action preconditions [36, 35], code execution errors [37], and syntactic errors in structured output [22]. Our work utilizes the same syntactic correction re-prompting technique as Skreta et al. [22], but we further introduce automatic detection and correction of semantic errors via re-prompting.
7 Conclusion and Limitations
@@
This paper presented AutoTAMP, a framework for using pre-trained LLMs as translators from language task descriptions to formal task specifications in STL via few-shot learning; a key contribution is that we show that our use of STL significantly improves the performance of the planner, and that the translation via automatic correction of syntax errors and a novel autoregressive re-prompting technique for semantic error checking and correction lead to further performance gains. Our experimental results show using LLMs to translate to task specifications that can be solved via a formal planner outperforms approaches that use LLMs directly as planners when handling tasks with complex geometric and temporal constraints.
Limitations 1) Though our results rely on using the best prompt out of several candidates, the prompt used may not be the optimal for eliciting the best performance of our LLMs. As a result, the individual results for each method may be improvable, but we expect the trends to persist and will continue to support the conclusion that LLMs are not well suited for direct complex TAMP. 2) Though the task success rates of AutoTAMP are good, the cost of the planning time is high, especially for multiple syntactic and semantic re-prompting iterations. Further work is needed that focuses on how to accelerate formal planners and decrease LLM inference latency. 3) Some tasks still fail due to translation errors. As part of future work, we aim to explore whether feedback via a human-in-the-loop framework can both improve translation and reduce AutoTAMP iterations.
Acknowledgments
We thank OpenAI for approving the usage of GPT-3/4 APIs.
The authors would also like to thank Charles Dawson, Oswin So, and Yue Meng for the early-stage discussion and final reviews of the work.
@@ Acknowledgments
This work was supported by ONR under Award N00014-22-1-2478 and MIT-IBM Watson AI Lab. However, this article solely reflects the opinions and conclusions of its authors.
Appendix A Recursive LLM Prompts
In this work, some methods need to re-prompt LLMs with context of the prompt history. Unlike the GPT-4 API that provides a template for this, the GPT-3 API template is for a single inference history; so, it is necessary to include any prompt history as context that indicates prior user input and LLM responses. We show an example in Figure 6.
Figure 6: Example of prompting GPT-3 API recursively. The prompt should indicate previous user prompts and LLM responses with tags (red).
Appendix B Convert Sub-task Sequence to STL
In order to evaluate LLM Task Planning methods, we use an STL planner [21] (this is the same planner we use for our AutoTAMP method) as the motion planner for generating trajectories. The LLM Task Planning methods generate output of the form [[$enter(room_{1})$, $t_{1}$], [$enter(room_{2})$, $t_{2}$],…, [$enter(room_{i})$, $t_{i}$], [$enter(room_{i+1})$, $t_{i+1}$]]. We automatically convert to STL to be consumed by the planner, i.e., finally[$t_{1}$, $t_{1}$] $enter(room_{1})$ and finally[$t_{2}$, $t_{2}$] $enter(room_{2})$ and finally[$t_{i}$, $t_{i}$] $enter(room_{i})$ and globally $not\_enter(walls)$. If the LLM Task Planning method also requires an executability check (e.g. SayCan, LLM Task Planing + Feedback), we apply the same conversion per sub-task; for example, given a sub-trajectory $i$, we would convert to STL of the form finally[$t_{i+1}-t_{i}$, $t_{i+1}-t_{i}$] $enter(room_{i+1})$ and globally $not\_enter(walls)$ to then be checked for executability.
Appendix C STL Syntax
Table 5: STL illustration
$\textbf{F}_{[a,b]}\phi$
True at time $t$ if there exists a time in the interval $[t+a,t+b]$ where $\phi$ is true.
$\phi\textbf{U}_{[a,b]}\varphi$
True at time $t$ if $\varphi$ is true for some time $t^{{}^{\prime}}$ in the interval $[t+a,t+b]$, and for all times between $t$ and $t^{{}^{\prime}}$, the formula $\phi$ holds.
$\textbf{G}_{[a,b]}\phi$
True at time $t$ if for all times in the interval $[t+a,t+b]$, the formula $\phi$ holds.
$\varphi\textbf{$\Rightarrow$}\textbf{F}_{[300,infinite]}\phi$
If $\varphi$ happens, then some time after the next 300 time steps $\phi$ should happen.
$\neg\varphi_{1}\textbf{$\lor$}(\varphi_{2}\textbf{U}_{[286,348]}\varphi_{3})$
$\varphi_{1}$ should not happen , or else $\varphi_{2}$ happens and continues until at some point during the 286 to 348 time units $\varphi_{3}$ happens .
Appendix D Pre-order and In-order STL Formats
As shown in Figure 7, each STL can be viewed as a binary tree with each node in the tree having at most two child nodes. When requesting NL-to-STL translation, the targeted tokens of STLs can be linearized in an in-order (left subtree, root, right subtree) or pre-order (root, left subtree, right subtree) manner. For this work, we prompt LLMs to generate STL in the pre-order format to avoid parentheses matching errors and for compatability with the STL planner. We note that the translation performance with GPT-4 is similar for both formats.
Figure 7: Illustration of pre-order and in-order formats of STL. (a) Different text representations of the formats for the same STL. (b) An example binary tree representation of STL.
Appendix E Manually Designed Instructions for HouseWorld Scenario
  1. 1.
    Navigate into the green bedroom and then go to the yellow restroom.
  2. 2.
    Travel to the yellow restroom, and then go to the pink bedroom. Visit both the kitchen and the dining room. Finally end in the red bathroom.
  3. 3.
    Navigate into the green bedroom and then go to the yellow restroom, but remember do not touch the livingroom at anytime.
  4. 4.
    Go to one room with cyan color, then enter the bedroom and stay there for 5 seconds, finally reach the restroom. remember always do not touch the two true blue areas.
  5. 5.
    Visit two rooms with color closest to red, but not the pure red color.
  6. 6.
    Visit every room that is not light pink or cyan.
  7. 7.
    Always go to the kitchen immediately before entering either bathroom. Go to the yellow room and the red room, but do not enter any blue rooms.
  8. 8.
    Move to the kitchen, then move to the bedroom without entering any blue rooms. Stay there for at least 10 seconds, then return to the kitchen via the living room.
  9. 9.
    Eventually end up in the master bedroom, but first go to both restrooms. After you have gone to each restroom, do not pass through any other rooms.
  10. 10.
    Visit every room at least once, but not go to the pink room until you are in a larger blue room.
Appendix F Keyframes of Trajectories for Multi-agent Scenarios.
Figure 8: Multi-agent task scenarios: Overcooked, Rover, and Wall. For Overcooked, the blue regions represent target ingredients, the center black square in an inadmissible region, and the yellow and green squares are agents. For Rover and Wall, the blue regions represent goals, the black lines represent walls, and the agents are the brown, yellow, orange, and green squares. In all scenarios, the agents occupy space and are consequently constrained by the inadmissible regions when moving through the environments.

### v3

@@ V RESULTS
TABLE I: Task success rates for single-agent scenarios. Each scenario’s constraints are listed in the table.
HouseWorld1 HouseWorld2 Chip’s Challenge
Soft Time Constraint Hard Time Constraint Hard Geometric Constraints
GPT-3 LLMs as Motion Planners End-to-end Motion Planning 0.0% 0.0% 0.0%
LLMs as Task Planners Task Planning (naive) 74.0% 36.0% 0.0%
SayCan 75.5% 36.0% 0.0%
Task Planning (feedback) 79.0% 40.0% 0.0%
LLMs as Translators No Corrections 28.0% 27.0% 29.0%
Syntax 49.0% 47.0% 66.0%
Syntax + Semantics (AutoTAMP) 62.0% 62.0% 74.3%
GPT-4 LLMs as Motion Planners End-to-end Motion Planning 9.5% 9.5% 0.0%
LLMs as Task Planners Task Planning (naive) 90.0% 45.0% 0.0%
Saycan 90.0% 47.5% 0.0%
Task Planning (feedback) 92.0% 49.0% 0.0%
LLMs as Translators No Corrections 43.5% 42.0% 42.7%
Syntax 59.5% 59.0% 70.0%
Syntax + Semantics (AutoTAMP) 82.5% 82.0% 87.7%
NL2TL + Syntax + Semantics - 83.5% 86.0%
TABLE II: Task success rates for multi-agent scenarios. Each situation has hard constraints on time and geometry.
@@
Hard Time & Geometric Constraints
GPT-3 LLMs as Motion Planners End-to-end Motion Planning 0.0% 0.0% 0.0%
LLMs as Task Planners Task Planning (naive) 13.3% 0.0% 7.0%
LLMs as Translators No Corrections 25.0% 22.0% 74.0%
Syntax Corrections 70.0% 35.0% 85.0%
Syntax + Semantic Corrections (AutoTAMP) 89.0% 60.7% 89.7%
GPT-4 LLMs as Motion Planners End-to-end Motion Planning 5.0% 0.0% 6.0%
LLMs as Task Planners Task Planning (naive) 17.0% 0.0% 47.0%
LLMs as Translators No Corrections 85.0% 46.0% 95.0%
Syntax Corrections 94.0% 67.0% 95.0%
Syntax + Semantic Corrections (AutoTAMP) 100.0% 79.0% 100.0%
NL2TL + Syntax + Semantic Corrections 100.0% 79.7% 100.0%
@@
We report the task success rates for the single-agent and multi-agent scenarios in Table I and Table II, respectively. For HouseWorld1 (Figure 4(a)) with no hard time constraint, we find that all methods using LLMs as task planners outperform our approach; whereas our approach can fail due to translation errors, this environment permits direct trajectories between any two positions and thus lacks geometric challenges that direct task planning methods will struggle with. When adding a strict time constraint (HouseWorld2), we see that such methods perform much worse while AutoTAMP’s success rate persists. For the other tasks that include geometric constraints, LLM End-to-end Motion Planning and Naive Task Planning both perform quite poorly. Unsurprisingly, we observe a general trend that GPT-4 outperforms GPT-3.
We find that most failures for LLM Task Planning methods result from task execution time violation and sequencing of actions for long-horizon tasks. For example, Chip’s Challenge requires the robot to efficiently collect keys for future doors. Also, the Naive Task Planning method fails to avoid collisions in the multi-agent scenarios. Failures for methods that translate to STL primarily are due to incorrect translation; while our re-prompting techniques help address this issue, there remain cases of poor translation.
Ablation Studies In Table I and Table II, we evaluate the impact of syntactic and semantic error correction on using LLMs to translate to STL. The results show that translation with no error correction has modest success across task scenarios, but both syntactic and semantic error correction significantly improve performance; this trend is present across all scenarios. We also evaluate replacing a pre-trained LLM for translation with a state-of-the-art modular translation pipeline, NL2TL, that uses a smaller LLM (T5-large) fine-tuned on a multi-domain corpus of 30K examples of instructions paired with their corresponding temporal logic expressions [21]; the error correction steps were still performed by GPT-4. Integrating NL2TL performs similarly to using a pre-trained LLM for translation, providing a modest improvement in HouseWorld2 and Rover. We note that incorporating the two re-prompting techniques for error correction is competitive with fine-tuning since we do not rely on additional data or training.
3D Simulation In supplemental videos, we demonstrate plans generated via AutoTAMP in two 3D simulated environments: a drone navigation scenario that requires reasoning about height, and a tabletop color sorting manipulation scenario. We did not incorporate the semantic check for these demos. The STL planner is directly applicable to the drone scenario using timed waypoints, as done in the 2D experiments. For manipulation tasks, we integrated a simple discrete planner to handle the dynamics mode transitions. We discuss this more in Section VII.
Physical Demonstrations We demonstrate AutoTAMP on physical differential-drive robots via the remotely-accessible Robotarium platform [25] for the Overcooked, Rover, Wall, and Chip’s Challenge scenarios. We track the planned trajectories using a low-level controller that also includes a control barrier function to prevent collisions between robots. This controller and the underlying nonlinear dynamics induce a tracking error; we account for this by padding obstacles at planning time. Obstacles are displayed in the robot workspace using an overhead projector. These physical demos provide evidence that our method can be applied to real-world navigation task and motion planning. They are included as part of supplemental videos.
VI RELATED WORK
Task and Motion Planning Planning for robotics involves both high-level, discrete planning of tasks [5] and low-level continuous planning of motions [26]; solving these simultaneously is referred to as task and motion planning [1]. Modern approaches either attempt to satisfy the motion constraints prior to action sequencing [27, 28, 29], find action sequences then satisfy the motion constraints [30, 31, 32, 19], or interleave these steps [33, 34, 35]. For tasks specified in temporal logic, existing methods either use multi-layer planning [36], like the aforementioned approaches, or direct optimization via a mixed-integer linear program [37, 23] or a non-linear program [38]. Our work focuses on translating natural language to STL, relying on [23] as a TAMP solver, but can be integrated with other STL-based planners.
@@ VI RELATED WORK
LLMs for TAMP Recent claims about the impressive reasoning capabilities of LLMs [6, 39] have led to interest in such models for task and motion planning. One approach is to directly use LLMs as planners [8, 9, 12, 11, 7, 10, 13]. Initial work showed that zero-shot generation of an action sequence from a high-level task description had relatively poor executability, but few-shot in-context learning, constraining output to admissible actions, and iterative action generation significantly improved performance [8]. Subsequent efforts grounded the primitive actions to motion control policies, using affordance functions to guide LLM-based task planning [9] and TAMP [12], also adding feedback [11]. Other work focused on how prompting can inform task execution[7, 13]. Despite these successes, however, there is evidence that LLMs perform poorly on more realistic tasks [15, 40], motivating different approaches. While we are interested in LLMs for TAMP, our work does not directly use LLMs as planners.
@@
Translating Language to Task Representations A natural alternative is to rely on dedicated planners by mapping from natural language to a planning representation. There is a rich history of parsing natural language into formal semantic representations [41, 42, 43], of which we only provide a relatively small sampling. The robotics community adopted parsing and other techniques to map language to such representations as lambda calculus [44, 45], motion planning constraints [46], linear temporal logic [47, 48, 49, 50], and signal temporal logic [51, 52], among others [53]. We refer readers to [54] for a more thorough review. To address challenges of data availability, task generalization, linguistic complexity, common sense reasoning, and more, recent work has applied LLMs to this translation problem. Modular approaches have used LLMs to extract referring expressions with corresponding logic propositions to then construct a full temporal logic specification [55, 21]. Relying on LLMs for direct translation, other work has mapped from language to PDDL goals [17] or full PDDL problems [56, 16]. Our work similarly translates to a task specification, but we can represent complex constraints (e.g. temporal), and we introduce a novel mechanism for automatic detection and correction of semantic errors. An interesting alternative maps language to code [57], which is highly expressive but does not easily optimize or provide behavior guarantees for long-horizon tasks.
@@
Re-prompting of LLMs The quality of LLM output is greatly improved with useful context, such as few-shot in-context learning for novel tasks [6]. LLMs for TAMP are typically also provided task-relevant information, such as environment state or admissible actions [10]. Re-prompting with additional context based on LLM output has been shown to be extremely beneficial, such as with iterative action generation [8], environmental feedback [11], inadmissible actions [8, 9, 12], unmet action preconditions [58, 56], code execution errors [59], and syntactic errors in structured output [20]. Our work uses the same syntactic correction re-prompting technique as [20], but we also introduce automatic detection and correction of semantic errors via re-prompting.
VII CONCLUSION
@@
This paper presented AutoTAMP, a framework for using pre-trained LLMs as both (1) translators from language task descriptions to formal task specifications (e.g. STL) via few-shot in-context learning and (2) checkers of syntactic and semantic errors via corrective re-prompting, in which we contributed a novel autoregressive re-prompting technique for semantic errors. Our experimental results show using LLMs to translate to task specifications that can be solved via a formal planner outperforms approaches that use LLMs directly as planners when handling tasks with complex geometric and temporal constraints.
We note a few limitations. First, though our results rely on using the best prompt out of several candidates, alternatives may elicit better performance. However, we expect the trends between methods to persist even with better prompts, supporting the conclusion that LLMs are not well suited for directly solving complex TAMP. Second, the cost of planning time is high, especially when there are multiple iterations of re-prompting. Further work is needed to address the runtime of formal planners and LLM inference. Third, the STL planner used in this work is not immediately applicable to manipulation tasks due to the optimization methods used in the planner; however, our approach does not depend on this specific planner, and we believe it can be integrated with STL planners more suitable for such TAMP domains.
VIII ACKNOWLEDGEMENTS
@@
This work was supported by ONR under Award N00014-22-1-2478, Army Research Laboratory under DCIST CRA W911NF-17-2-0181, and MIT-IBM Watson AI Lab. This article solely reflects the conclusions of its authors.